# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ 5.6 Uncertainty handling
Given the uncertainty inherent in health decision-making as well as the high likelihood of ambiguity, noise, and missing values in sensor observations, health monitoring systems are greatly enhanced by being able to handle uncertainty. Despite this, only 19 of the systems addressed some aspect of this. The approaches used to handle uncertainty are summarised in Table 7 and discussed in detail in the remainder of this subsection.
@@
Replacing or eliminating missing or invalid sensor data Ali et al. [77, 78]; Hooda and Rani [91]; Hussain and Park [93]; Reda at al. [73]; Rhayem et al. [105]; Titi et al. [107]
@@ 5.6.2 Bayesian networks
Bayesian networks are well known for managing uncertainty and have been widely used in the health domain [144]. Kordestani et al. [98] use a Bayesian network for probabilistic diagnosis of acute kidney injury. The Bayesian network models immediate (short-term) and background (long-term) causes of acute kidney injury, as well as its symptoms. They used experts to determine the conditional probabilities of the presence of acute kidney injury given these variables. Similarly, Mcheick et al. [101] represent risk factors for stroke using a Bayesian network.
@@ 5.6.4 Preprocessing sensor data
Uncertainty can stem from various factors in sensor data, including ambiguous or imprecise readings, noise, or missing values caused by sensor malfunctions or network failures [21, 20]. Several systems have addressed the issue of missing values in sensor data. Ali et al. [77, 78] replace them with mean and median values from existing data, while Hooda and Rani [91] replace them using the preceeding value. Rhayem et al. [105] take the approach of removing any missing or unusual values, for example those outside the device measurement ranges. Similarly, Titi et al. [107] and Reda et al. [73] use rules to check whether sensor data falls within the expected minimum and maximum bounds. In their proposed system, Hussain and Park [93] use the Pan-Tompkins algorithm to detect the QRS complex in the ECG. This identifies beats without a QRS complex, which may be premature, missing, or ectopic, and are subsequently eliminated. To deal with noisy data, a few systems use filters to improve signal quality. Ali et al. [77, 78] use a Kalman filter to remove noise, while Garcia-Valverde et al. [88] use a moving average filter for the same purpose.
@@ 5.6.5 The role of Semantic Web technologies
Semantic Web technologies provide limited inherent support for uncertainty handling, but this can be mitigated through the techniques discussed in this section. Several extensions for Semantic Web standards have been proposed in the literature that make use of fuzzy logic and Bayesian inference, such as BayesOWL [153] and Bayes-SWRL [154], probabilistic extensions for OWL and SWRL respectively, as well as fuzzyDL [155], a fuzzy ontology reasoner. However, none of the selected systems that implement uncertainty handling techniques report using any such extensions, opting instead to define custom solutions. Another way that Semantic Web technologies can support uncertainty handling is through using ontology reasoners for inconsistency detection. Missing values or otherwise invalid data can be detected through reasoners such as HermiT [156] and Pellet [157] based on specified and inferred axioms; the use of these tools are discussed in greater detail in Section 6. Additionally, rule-based reasoning can be combined with ontologies to detect invalid data, an approach used by a few of the selected systems.
@@ 5.7.4 Ethics and regulatory compliance
The high-stakes nature of the health domain necessitates careful consideration of ethical issues. Although there are many benefits of technology-enabled personal health monitoring, there are also potential harms that it exposes. Many of these ethical issues overlap with the challenges already discussed, such as situation analysis (how accurate and reliable are the detected and predicted situations?), decision support (how appropriate are the suggested recommendations and how much autonomy do system users have?), explainability (to what extent can system outputs be understood?), and security and privacy (how secure is user data and how is consent managed?). An additional ethical concern is the cascade of care, a phenomenon in which incidental findings from screenings or monitoring result in further clinical care. Some of these ethical considerations can be enforced through regulation. Readers seeking further exploration on this challenge may consult the following articles: Morley et al. [167] comprehensively map the ethics of AI in healthcare; Kwan et al. [168] highlight the ethical issues associated with health monitoring applications; Nittari et al. [169] review ethical and legal challenges in telemedicine more broadly; Hassanaly and Dufour [170] explore the regulation of mobile health applications in the United States, the European Union, and France; and Thapa and Camtepe [159] explore legality and regulatory compliance from the perspective of security and privacy.
5.8 Summary
This section has provided an in-depth analysis of seven key challenges that must be addressed in health monitoring systems, as well as a brief discussion of additional challenges that are important in such systems. The role played by Semantic Web technologies in overcoming the seven key challenges has been critically examined. Additionally, non-semantic techniques that are incorporated in the systems have also been discussed. A full list of reused semantic resources, systems that reuse them, and the challenges they address can be found in Table 16 in the appendix. These resources include ontologies, knowledge graphs, and linked data, as well as vocabularies, taxonomies, and classifications. To summarise the section, we conduct an assessment of each system based on the seven key challenges. Finally, we discuss the results of the assessment, highlighting the challenges that are most neglected among the systems.
5.8.1 Challenges assessment
Table 8 summarises the different aspects related to the seven key challenges. While we consider these aspects to be highly important for achieving effective sensor-based personal health monitoring, it is possible that some of them may exceed the requirements for specific health conditions and may therefore not be essential in those particular cases. To assess the degree to which each system tackles the seven challenges, we use a rating scheme based on the identified aspects. A score of 1 point is assigned to each system for every aspect that is addressed. We use a five-point rating scale as follows:
@@ 5.8.1 Challenges assessment
    High: between 70% and 90% of the aspects are addressed
  5. 5.
    Very High: more than 90% of the aspects are addressed
@@
Explainability 1. The system uses inherently interpretable techniques as part of the situation analysis process (for example when classifying sensor data), or else applies a post hoc explainability method. 2. The system uses inherently interpretable techniques as part of the decision support process (for example when making recommendations), or else applies a post hoc explainability method.
Table continued on next page.
Important aspects related to the seven key challenges.
Challenge Aspects
3. It is mentioned or demonstrated that the system presents explicit explanations to the user for generated situations. 4. It is mentioned or demonstrated that the system presents explicit explanations to the user for generated decisions. 5. The presented explanations meet two or more of the criteria for good explanations as explained by Molnar [140], e.g. contrastiveness, conciseness, generalisability, and a focus on abnormalities. 6. The presented explanations are tailored to different target audiences (monitored individuals, clinicians, caregivers, regulators, etc), e.g. by providing different levels of detail or highlighting different aspects of the situation or decision.
Uncertainty handling 1. The system is able to handle uncertainty in the situation analysis process. 2. The system is able to handle uncertainty in the decision support process. 3. The system is able to handle missing, noisy, or otherwise invalid sensor data.
5.8.2 Discussion
Table 9 shows the number of systems with a particular rating for each challenge. Note that because situation detection, situation prediction, and uncertainty handling each have three aspects as per Table 8, a rating of “high” is not applicable based on the five-point rating scale. The combined radar chart in Figure 4 provides a visualisation of how well all the systems address the seven challenges. Separate radar charts for the individual systems are also available [^30] https://public.flourish.studio/visualisation/17843904, and the individual system ratings are shown in Table 17 in the appendix.
It is evident that more work is needed to address situation prediction, explainability, uncertainty handling, and to a lesser extent, interoperability and decision support. Situation prediction stands out as the most neglected challenge, with 30 of the 43 systems failing to address it altogether. This is closely followed by uncertainty handling, which 23 systems do not address. It is also notable that none of the selected systems have achieved a rating higher than medium for situation prediction, explainability, and uncertainty handling. Interoperability is not very well addressed among the systems, with most scoring a low rating for this challenge, although two of the systems achieve a high score. Moreover, although seven of the systems score either high or very high on decision support, most only moderately address it, and three do not address it at all. While nine systems score very highly on situation detection, the remainder score either a low or medium rating. However, all of the selected systems consider this challenge, making it generally well addressed. The same is true of context awareness, with 22 systems achieving a medium score, 17 achieving a high score, and one achieving a very high score.
@@ 5.8.2 Discussion
Rating Challenge Interoperability Context awareness Situation detection Situation prediction Decision support Explainability Uncertainty handling
3 0 0 30 3 0 23
Low 25 3 12 12 19 35 15
Medium 13 22 22 1 14 8 5
High 2 17 N/A N/A 5 0 N/A
Very High 0 1 9 0 2 0 0
@@ 6.1 Data and devices
The data collection methodology varies among the systems. Many systems used existing datasets from publicly available repositories such as PhysioNet [^31] https://physionet.org and the University of California, Irvine (UCI) ML repository [^32] https://archive.ics.uci.edu. These systems and the datasets they use are summarised in Table 10.
@@
Ali et al. [77] Pima Indians diabetes dataset Multiparameter Intelligent Monitoring in Intensive Care (MIMIC-II) Drug review dataset UCI ML Repository PhysioNet UCI ML Repository
@@
Hooda and Rani [91] Pima Indians diabetes dataset Heart disease dataset (Cleveland) UCI ML Repository UCI ML Repository
Table continued on next page.
Existing health datasets used.
[…]
Kordestani et al. [98] Chronic kidney disease dataset Dermatology dataset UCI ML Repository UCI ML Repository
Mavropoulos et al. [100] Heterogeneity Human Activity Recognition dataset CoNNL2003 dataset UCI ML Repository Sang et al. [172]
Peral et al. [104] Diabetes dataset Health Facts database UCI ML Repository Strack et al. [173]
@@
A smaller number of the systems used data collected from participants rather than existing data. For example, Ali et al. [79] collected data from 44 diabetes patients, while Esposito et al. [86] collected data from 10 healthy volunteers. Other systems that used this approach are those proposed by De Brouwer et al. [83], Hristoskova et al. [92], Hussain and Park [93], Stavropoulos et al. [74], Vadillo et al. [108], and Villarreal et al. [109]. Another approach was to simulate or manually generate the data. This was done by Alti et al. [80] who simulated temperature and camera data; Chatterjee et al. [81], who simulated the sensor, interview, and questionnaire data of four dummy participants; and Zafeiropoulos et al. [75], who simulated the sensor observations and health records of three virtual patients. Mchiek et al. [101] similarly generated 513 data records. Additionally, Stavropoulos et al. [74] simulated records in order to test the scalability of their proposed system. A significant number of the systems indicated the types of data and sensors supported by the systems, but did not mention the source of the data. It is unclear whether these systems were validated using actual sensor data, beyond a theoretical validation of the system functionality.
Less than half of the systems gave specific details of the devices used for data collection. While most of these were commercially available devices, the systems proposed by Kim et al. [97] and Lopes de Souza et al. [97] used custom-made prototypes. Table 11 indicates the types and descriptions of devices mentioned.
@@
Esposito et al. [86] Commercial Amiigo wristband Omron HJ-112 digital pocket pedometer BT, HR, SpO2 Acceleration
Hristoskova et al. [92] Commercial A&D UA-767PBT blood pressure monitor Nonin Avant 4000 digital pulse oximeter A&D UC-321PBT weight scale Welch Allyn Cardio Perfect 12 Lead ECG BP, HR SpO2 Body weight ECG
Hussain and Park [93] Commercial BioNomadix respiration (RSP) with ECG amplifier ECG patch from Life Science Technology Inc. ECG
[…]
@@
Table continued on next page.
Devices mentioned for data collection in the selected systems.
System Device type Name/description Data type
@@
Mavropoulos et al. [100] Commercial Samsung Galaxy S3 Mini smartphone LG Nexus 4 smartphone Video
Reda et al. [73] Commercial Fitbit Jawbone Step count, HR, calories burned
Spoladore et al. [106] Commercial Polar H1 chest strap COSMED E100 cycle ergometer ECG
Stavropoulos et al. [74] Commercial Fitbit Charge 3 Step count, sleep stages, HR
Vadillo et al. [108] Commercial Arduino e-Health Sensor Platform Tunstall Lifeline Connect+ home unit BG, HR, BT, BP, SpO2 Motion

### v3

@@ 5.7 Uncertainty handling
Given the uncertainty inherent in health decision-making as well as the high likelihood of ambiguity, noise, and missing values in sensor observations, health monitoring systems are greatly enhanced by being able to handle uncertainty. Despite this, only 22 of the systems addressed some aspect of this [109, 110, 111, 112, 88, 90, 117, 91, 92, 94, 119, 120, 97, 123, 125, 127, 128, 129, 106, 83, 100, 103]. The approaches used to handle uncertainty are summarised in Table 8 and discussed in detail in the remainder of this subsection.
@@
Replacing or eliminating missing or invalid sensor data Ali et al. [110, 111]; Garcia-Moreno et al. [91]; Hooda and Rani [94]; Hussain and Park [120]; Martella et al. [125]; Reda at al. [106]; Rhayem et al. [83]; Titi et al. [100]
@@ 5.7.2 Bayesian networks
Bayesian networks are well known for modelling uncertainty and have been widely used in the health domain [168]. Kordestani et al. [123] use a Bayesian network for probabilistic diagnosis of acute kidney injury. The Bayesian network models immediate (short-term) and background (long-term) causes of acute kidney injury, as well as its symptoms. They used experts to determine the conditional probabilities of the presence of acute kidney injury given these variables. Similarly, Mcheick et al. [127] represent risk factors for stroke using a Bayesian network.
@@ 5.7.3 Nonmonotonic reasoning
5.7.4 Probabilistic risk classification
Hristoskova et al. [119] account for uncertainty in situation prediction by classifying patients into risk stages based on the four-year probability of congestive heart failure. They use probabilistic rules defined using SWRL to analyse risk factors and classify individuals based on the Framingham Risk Score.
@@ 5.7.5 Preprocessing sensor data
Uncertainty can stem from various factors in sensor data, including ambiguous or imprecise readings, noise, or missing values caused by sensor malfunctions or network failures [31, 30]. 16 systems have addressed the issue of missing or invalid values in sensor data. Ali et al. [110, 111] replace them with mean and median values from existing data, while Hooda and Rani [94] replace them using the preceding value. Rhayem et al. [83] take the approach of removing any missing or unusual values, for example those outside the device measurement ranges. Similarly, Kilintzis et al. [97], Martella et al. [125], Titi et al. [100] and Reda et al. [106] use rules to check whether sensor data falls within the expected minimum and maximum bounds. In their proposed system, Hussain and Park [120] use the Pan-Tompkins algorithm to detect the QRS complex in the ECG. This identifies beats without a QRS complex, which may be premature, missing, or ectopic, and are subsequently eliminated. While Garcia-Moreno et al. [91] mention missing value imputation as part of their ML pipeline, they do not specify a technique for this. To deal with noisy data, three systems use filters to improve signal quality. Ali et al. [110, 111] use a Kalman filter to remove noise, while Garcia-Valverde et al. [92] use a moving average filter for the same purpose.
@@ 5.7.6 The role of Semantic Web technologies
As discussed in the previous challenges, Semantic Web technologies provide limited inherent support for uncertainty handling, but this can be mitigated through the techniques discussed in this section. Several extensions for Semantic Web standards have been proposed in the literature that make use of fuzzy logic and Bayesian inference, such as BayesOWL [184] and Bayes-SWRL [185], probabilistic extensions for OWL and SWRL respectively, as well as fuzzyDL [150], a fuzzy ontology reasoner. However, none of the selected systems report using any such extensions, opting instead to define custom solutions. Another way that Semantic Web technologies can support uncertainty handling is through using ontology reasoners for inconsistency detection. Missing values or otherwise invalid data can be detected through ontology reasoners based on specified and inferred axioms; the use of these tools are discussed in greater detail in Section 6. Additionally, rule-based reasoning can be combined with ontologies to detect invalid data, an approach used by five of the selected systems [83, 97, 125, 100, 106]. Finally, Semantic Web technologies can indirectly support uncertainty handling by representing data quality properties that affect uncertainty. For instance, Garcia-Moreno et al. [91] model quality properties of sensor data using an ontology, including correctness and completeness.
@@ 5.8.4 Ethics and regulatory compliance
The high-stakes nature of the health domain necessitates careful consideration of ethical issues. Although there are many benefits of technology-enabled personal health monitoring, there are also potential harms that it exposes. Many of these ethical issues overlap with the challenges already discussed, such as situation detection and situation prediction (how accurate and reliable are the detected and predicted situations?), decision support (how appropriate are the suggested recommendations and how much autonomy do system users have?), explainability (to what extent can system outputs be understood?), and security and privacy (how secure is user data and how is consent managed?). An additional ethical concern is the cascade of care, a phenomenon in which incidental findings from screenings or monitoring result in further clinical care. This may cause undue anxiety and psychological harm to the monitored individual, while also potentially leading to costly or invasive diagnostic procedures. Some of these ethical considerations can be enforced through regulation. Readers seeking further exploration on this challenge may consult the following articles: Morley et al. [195] comprehensively map the ethics of AI in healthcare; Kwan et al. [196] highlight the ethical issues associated with health monitoring applications; Nittari et al. [197] review ethical and legal challenges in telemedicine more broadly; Hassanaly and Dufour [198] explore the regulation of mobile health applications in the United States, the European Union, and France; and Thapa and Camtepe [187] explore legality and regulatory compliance from the perspective of security and privacy.
5.9 Challenges assessment
In this subsection, we conduct an assessment of each system based on the seven key challenges. Table LABEL:challenges-aspects summarises the different aspects related to the challenges. While we consider these aspects to be highly important for achieving effective sensor-based personal health monitoring, it is possible that some of them may exceed the requirements for specific health conditions or applications, and may therefore not be essential in those particular cases. To assess the degree to which each system tackles the seven challenges, we use a rating scheme based on the identified aspects. A score of 1 point is assigned to each system for every aspect that is addressed. All challenges are equally weighted. We use a four-point rating scale as follows:
@@ 5.9 Challenges assessment
    High: more than 70% of the aspects are addressed
@@
Explainability 1. The system uses inherently interpretable techniques as part of the situation detection or situation prediction processes (for example when classifying sensor data), or else applies a post hoc explainability method. 2. The system uses inherently interpretable techniques as part of the decision support process (for example when making recommendations), or else applies a post hoc explainability method. 3. It is mentioned or demonstrated that the system presents explicit explanations to the user for generated situations. 4. It is mentioned or demonstrated that the system presents explicit explanations to the user for generated decisions. 5. The presented explanations meet two or more of the criteria for good explanations as explained by Molnar [165], e.g. contrastiveness, conciseness, generalisability, and a focus on abnormalities. 6. The presented explanations are tailored to different target audiences (monitored individuals, clinicians, caregivers, regulators, etc), e.g. by providing different levels of detail or highlighting different aspects of the situation or decision.
Uncertainty handling 1. The system is able to handle uncertainty in the situation detection or situation prediction processes. 2. The system is able to handle uncertainty in the decision support process. 3. The system is able to handle missing, noisy, or otherwise invalid sensor data.
Table 10 shows the number of systems with a particular rating for each challenge. The combined radar chart in Figure 4 provides a visualisation of how well all the systems address the seven challenges. Separate radar charts for the individual systems are also available [^33] https://public.flourish.studio/visualisation/17843904, and the individual system ratings are shown in Table LABEL:challenges-summary-systems in the appendix.
@@
Rating Challenge Interoperability Situation detection Situation prediction Decision support Context awareness Explainability Uncertainty handling
3 0 36 3 0 0 26
Low 24 15 11 22 3 40 17
Medium 19 24 1 20 23 8 5
High 2 9 0 3 22 0 0
@@
5.10 Summary
This section has provided an in-depth analysis of seven key challenges that must be addressed in health monitoring systems, as well as a brief discussion of additional challenges that are important in such systems. The role played by Semantic Web technologies in overcoming the seven key challenges has been critically examined. Additionally, non-semantic techniques that are incorporated in the systems have also been discussed. A full list of reused semantic resources, systems that reuse them, and the challenges they address can be found in Table LABEL:reused-summary in the appendix. These resources include ontologies, knowledge graphs, and linked data, as well as vocabularies, taxonomies, and classifications. To summarise the section, we discuss the results of the assessment, highlighting the challenges that are most neglected among the systems.
Based on the challenges assessment described in Section 5.9, it is evident that more work is needed to address situation prediction, uncertainty handling, explainability, and to a lesser extent, interoperability and decision support. Situation prediction stands out as the most neglected challenge, with 36 of the 48 systems failing to address it altogether. This is closely followed by uncertainty handling, which 26 systems do not address at all. Although all of the selected systems have some degree of explainability through their use of inherently interpretable Semantic Web technologies, 40 of them have a low rating for this challenge, with the remaining eight having a medium rating. Notably, none of the selected systems achieve a high rating for situation prediction, uncertainty handling, or explainability. Interoperability is not very well addressed among the systems, with half of them (24) scoring a low rating for this challenge, although two systems achieve a high score. Similarly, only three systems achieve a high score for decision support. There is a nearly even split between low and medium ratings, with 22 and 20 systems respectively, while three systems do not provide decision support at all. Context awareness is the best addressed challenge among the systems, with 22 achieving a high score, 23 achieving a medium score, and the remaining three attaining a low score. Situation detection is also fairly well addressed, with nine high scores, 24 medium scores, and 15 low scores.
@@ 6.1 Data and devices
The data collection methodology varies among the systems. 14 systems used existing datasets from publicly available repositories such as PhysioNet [^34] https://physionet.org and the University of California, Irvine (UCI) ML repository [^35] https://archive.ics.uci.edu. These systems and the datasets they use are summarised in Table LABEL:existing-datasets.
@@
Ali et al. [110] 1. Pima Indians diabetes dataset 1. UCI ML Repository
2. MIMIC-II 2. PhysioNet
3. Drug review dataset 3. UCI ML Repository
@@
Hooda and Rani [94] 1. Pima Indians diabetes dataset 1. UCI ML Repository
2. Heart disease dataset (Cleveland) 2. UCI ML Repository
Ivașcu and Negru [95] Mobile health dataset UCI ML Repository
Kordestani et al. [123] 1. Chronic kidney disease dataset 1. UCI ML Repository
2. Dermatology dataset 2. UCI ML Repository
1. Heterogeneity Human Activity Recognition dataset 1. UCI ML Repository
Mavropoulos et al. [126] 2. CoNNL2003 dataset 2. Sang et al. [200]
Peral et al. [130] 1. Diabetes dataset 1. UCI ML Repository
2. Health Facts database 2. Strack et al. [201]
@@
10 systems [112, 115, 90, 97, 91, 119, 120, 99, 101, 102] used data collected from participants rather than existing data. For example, Ali et al. [112] collected data from 44 diabetes patients, while Esposito et al. [90] collected data from 10 healthy volunteers. Another approach was to simulate or manually generate the data. This was done by Alti et al. [113] who simulated temperature and camera data; Bampi et al. [87] who simulated time-varying sensor data; Chatterjee et al. [114], who simulated the sensor, interview, and questionnaire data of four dummy participants; Martella et al. [125], who used a script to simulate sensor-generated data streams; and Zafeiropoulos et al. [107], who simulated the sensor observations and health records of three virtual patients. Mcheik et al. [127] similarly generated 513 data records. Additionally, Stavropoulos et al. [99] simulated records in order to test the scalability of their proposed system. 18 systems [109, 88, 116, 89, 117, 96, 121, 124, 128, 129, 106, 98, 100, 101, 105, 131, 133, 103] indicated the types of data and sensors supported by the systems, but did not mention the source of the data. It is unclear whether these systems were validated using actual sensor data, beyond a theoretical validation of the system functionality. Only 19 of the systems gave specific details of the devices used for data collection. Table LABEL:devices-used indicates these systems and the types and descriptions of devices mentioned. While most of these are commercially available devices, the systems proposed by Kim et al. [122] and Lopes de Souza et al. [122] used custom-made prototypes.
@@
1. Amiigo wristband 1. BT, HR, SpO2
Esposito et al. [90] Commercial 2. Omron HJ-112 digital pocket pedometer 2. Acceleration
Garcia-Moreno et al. [91] Commercial 1. Samsung Gear S3 smartwatch 1. HR
2. Empatica E4 wristband 2. Accelerometer, electrodermal activity, HR, BT
1. A&D UA-767PBT blood pressure monitor 1. BP, HR
2. Nonin Avant 4000 digital pulse oximeter 2. SpO2
3. A&D UC-321PBT weight scale 3. Body weight
Hristoskova et al. [119] Commercial 4. Welch Allyn Cardio Perfect 12 Lead ECG 4. ECG
Hussain and Park [120] Commercial 1. BioNomadix respiration (RSP) with ECG amplifier ECG
2. ECG patch from Life Science Technology Inc.
Ivașcu and Negru [95] Commercial Shimmer2 Acceleration, ECG
@@
Martella et al. [125] Commercial 1. Fitbit wristband 1. HR, BT
2. Aircare device 2. TVOC, CO2, particulate matter, sound pressure level, temperature, relative humidity, ambient light, atmospheric pressure
1. Samsung Galaxy S3 Mini smartphone
Mavropoulos et al. [126] Commercial 2. LG Nexus 4 smartphone Video
Reda et al. [106] Commercial 1. Fitbit wristband
2. Jawbone wristband Step count, HR, calories burned
1. Polar H1 chest strap
Spoladore et al. [98] Commercial 2. COSMED E100 cycle ergometer ECG
Stavropoulos et al. [99] Commercial Fitbit Charge 3 wristband Step count, sleep stages, HR
1. Arduino e-Health sensor platform 1. BG, HR, BT, BP, SpO2
Vadillo et al. [101] Commercial 2. Tunstall Lifeline Connect+ home unit 2. Motion